OpenStudy (anonymous):

True or false? The inverse of the function y = 2x + 8 is y = x – 8/ 2

OpenStudy (hoblos):

true

OpenStudy (radar):

y=2x+8 exchange x and y getting x=2y+8 2y=x-8 y= x/2-4 Looks like answer should be false

OpenStudy (radar):

To be sure I am going to try and verify my answer \[f(f ^{-1}(x)=f((x/2)-4)=2((x/2)-4)+8=x-8+8=x\] Seems to check out.

OpenStudy (hoblos):

but (x – 8)/ 2 = (x/2) - 4 so where is the wrong part!!

OpenStudy (radar):

Well I see x - 8/2 not (x-8)/2 Are you assuming that is what it means??

OpenStudy (hoblos):

yes i think x - 8/2 means (x-8)/2 @FC Barcelona13 check it... if it is (x-8)/2 then the answer is TRUE if not then the answer is FALSE

OpenStudy (radar):

that should remove all doubt hoblos

OpenStudy (radar):

It would be nice if all ambiguity was eliminated in the posted questions. A prudent use of parenthesis could help a lot.